OpenStudy (anonymous):

Find the derivative of the function y=ln (1+e^x/ 1-e^x)

OpenStudy (anonymous):

can we use logarithmic properties here?

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

yes, but first we can break it up and make it easier on ourself

OpenStudy (anonymous):

ya.it would be better ....using log both side

OpenStudy (anonymous):

2e^x/(1-e^2x)

OpenStudy (anonymous):

yes how did you get that

OpenStudy (anonymous):

yeah thats the answer but to get there we can write it as: y=ln(1+e^x)-ln(1-e^x)

OpenStudy (anonymous):

Then we differntiate both sides

myininaya (myininaya):

\[y=\ln(1+e^x)-\ln|1-e^x|=>y'=\frac{(1+e^x)'}{1+e^x}-\frac{(1-e^x)'}{1-e^x}\]

OpenStudy (anonymous):

y=sin\[\sqrt[3]{x}\] +\[\ \sqrt[3]{sinx}\]